OpenStudy (anonymous):

|x|+|x-2|+|x-4|<5 x between?

OpenStudy (anonymous):

the answer is 1<x<3 but how they made it?

OpenStudy (gorv):

|x|= x for x>=0 -x for x<0 |x-2|=x-2 for x >=2 -(x-2) for x<2 |x-4|=x-4 for x>=4 -(x-4) for x<4 so for x>4 x+x-2+x-4=3x-6<5 ==> 3x<5+6 x<11/3

OpenStudy (gorv):

for x<0

OpenStudy (gorv):

-x-x+2-x+4<5 -3x+6<5 -3x<-1 x<1/3